### v1

# Holstein polarons, Rashba-like spin splitting and Ising superconductivity in electron-doped MoSe2

Abstract
Interaction between electrons and phonons in solids is a key effect defining physical properties of materials such as electrical and thermal conductivity. In transitional metal dichalcogenides (TMDCs) the electron-phonon coupling results in the creation of polarons, quasiparticles that manifest themselves as discrete features in the electronic spectral function. In this study, we report the formation of polarons at the alkali-dosed MoSe2 surface, where Rashba-like spin splitting of the conduction band states is caused by an inversion-symmetry breaking electric field. In addition, we observe the crossover from phonon-like to plasmon-like polaronic spectral features at MoSe2 surface with increasing doping. Our findings support the concept of electron-phonon coupling mediated superconductivity in electron-doped layered TMDC materials, observed using ionic liquid gating technology. Furthermore, the discovered spin-splitting at the Fermi level could offer crucial experimental validation for theoretical models of Ising-type superconductivity in these materials.
dichalcogenides, polarons, superconductivity
I Introduction
Polarons are collective excitation of electrons interacting with the lattice. These quasiparticles describe an electron moving in a crystal with positively charged ions displacing from their equilibrium positions to effectively screen negative electron charge Landau (1933); Landau and Pekar (1948). This coupling between the electron and lattice vibrations (phonon cloud) lowers the electron mobility and increases the electron’s effective mass.
The resulting polaronic bound state can be described by Fröhlich or Holstein Hamiltonians. The Fröhlich model describes the interaction of electrons with optical phonons, which leads to the formation of large polarons, whose radii are much larger than the lattice constant. The Holstein model, on the other hand, describes the interaction with acoustic phonons, which leads to the formation of small polarons, whose radii are of the same order of magnitude as the lattice constant Emin (2013).
Formation of the polarons due to a momentum dependent electron-phonon coupling leads to new features in the electronic spectral function at specific electron valleys, e.g. the emergence of discrete energy levels and discontinuities in the density of states Franchini et al. (2021). In a recent study of alkali-metal dosing on MoS2 surface such discontinuities in density of states have been experimentally observed at the ${\rm K}$ points of the Brillouin zone using angle-resolved photoemission spectroscopy (ARPES) Kang et al. (2018). Understanding the changing nature of the quasiparticles as a function of doping is an important prerequisite for understanding the onset of superconductivity observed in highly electron-doped MoS2 under liquid ionic gating Ye et al. (2012); Lu et al. (2015).
Theoretical considerations point to two key novelties in the case of multi-valley electron-doped TDMCs. First, polaron modes should be observed around both ${\rm K}$ and $\Sigma$ valleys in MoS2 Garcia-Goiricelaya et al. (2019).
Figure 1: Phonon-mediated polarons in electron doped MoSe2: (a) Multi-valley scattering of conduction electrons in the first Brillouin zone of MoSe2; (b) Electron-like parabolic band at ${\rm K}$-valley with Holstein polarons simulated by momentum-average model Goodvin et al. (2006).
Figure 1(a) shows the multi-valley scattering between $\Sigma$ and ${\rm K}$ valleys in the Brillouin zone. The electron scattering vectors $q_{1}$ and $q_{2}$ correspond to either $\Gamma$-${\rm K}$ or $\Gamma$-${\rm M}$ phonon wave vectors. For both MoS2 and MoSe2 the acoustic phonon modes have similar energies for ${\rm K}$ and ${\rm M}$ and the dominant contribution to electron-phonon coupling arises from the highest energy acoustic phonon mode. This behaviour has been demonstrated both experimentally Zeng et al. (2012) and theoretically Kaasbjerg et al. (2012) for MoS2. However, only ${\rm K}$ valley Holstein polarons (Fig. 1(b)) have been observed in the ARPES study of Rb-dosed MoS2 surface  Kang et al. (2018). Second, the spin-orbit coupling (SOC) should induce the splitting of the electron bands at $\Sigma$ and $\Sigma$’ valleys due to inversion symmetry breaking on the surface with alkali metal dosing. However, in all previous ARPES studies of alkali-dosed semiconducting TMDCs the spin-split features at $\Sigma$ valley have not been resolved Zhang et al. (2014); Eknapakul et al. (2014); Alidoust et al. (2014); Riley et al. (2015); Miwa et al. (2015); Zhang et al. (2016); Kim et al. (2016, 2017); Kang et al. (2017, 2018); Han et al. (2021).
Figure 2: Evolution of the MoSe2 electronic structure with Rb dosing: (a-a5) ARPES spectra of $\Gamma$-${\rm K}$ conduction band dispersions measured near the Fermi energy for dosing 0, 20, 60, 100, 220, 580s. (b-b5) Corresponding valence band spectra in a wide energy range.
The isostructural MoSe2 compound is the perfect model system for probing the multi-valley physics induced by electron-doping. Density functional theory (DFT) calculations show that in contrast to MoS2, the conduction band minimum (CBM) in bulk MoSe2 is at the $\Sigma$, rather than at ${\rm K}$ valley (see Supplementary Fig. S1). For the alkali-metal dosing on MoSe2 surface one expects to have larger SOC band-splitting as in the case of MoS2. The accessibility of the $\Sigma$ valley, as well as the ${\rm K}$ valley expected at higher doping Kim et al. (2016, 2017), makes MoSe2 the best candidate to study physics of multi-valley polaron modes using high-resolution ARPES.
II Results and Discussion
II.1 Evolution of the electronic band structure with dosing
In our ARPES data shown in Fig. 2 the spectral intensity appears at the Fermi level, marking the occupation of the conduction band and corresponding semiconductor to conductor transition with the Rb dosing on the MoSe2 surface. After initial dosing the maximum spectral weight at the Fermi level is observed around $\Sigma$ valley, showing an indirect band gap $E_{g}^{\Gamma-\Sigma}$ $\sim$1.4 eV. However, with increasing amount of dosed Rb the relative intensity of $\Sigma$ valley decreases, whereas the intensity of ${\rm K}$ valley increases. Correspondingly, it appears that the conduction band minimum shifts from $\Sigma$ to ${\rm K}$ valley as in case of monolayer MoSe2 (see Fig. 2(a1-a3)). Withing observed energy broadening the band minima at $\Sigma$ and ${\rm K}$ are very close (see Supplementary Fig. S5), incidentally making electron doped MoSe2 perfect system for studying multi-valley physics. Simultaneously with the shift of the CBM intensity, the intensity of the valence band compared to the conduction band decreases (see Fig. 2 (b1-b3)). This first crossover at around 40 s of Rb dosing marks the transition from the electron doped bulk to the more “monolayer-like” doped surface conduction band structure Eknapakul et al. (2014); Riley et al. (2015); Kang et al. (2017); Kim et al. (2017).
The accumulation of alkali atoms on the surface and increased electron doping of the top-layer causes a corresponding increase of the electric field, and results in the amplification of the inversion symmetry-breaking at the surface. At dosing above this first crossover the measured electronic structure is a superposition of the bands from both the bulk material and the doped surface. This observation is supported by our Se 3d core level and ${\rm K}$-valley valence band data (see Supplementary Fig. S4 and Fig. S5 correspondingly).
At Rb dosing times above 120 s the relative intensity of the valence bands increases due to the interlayer intercalation of Rb atoms Eknapakul et al. (2014). At the same time the energy difference between bulk and surface Se 3d core levels or “effective chemical shift” saturates and does not increase with Rb deposition time (see Supplementary Fig. S4). This corresponds to the second crossover from surface doping to interlayer intercalation regime. For dosing times above 340 s one start to observe build up of the photoemission intensity at the Fermi level around $\Gamma$-point as seen in Fig. 2(a5). This corresponds to the formation of the metallic Rb states on the MoSe2 surface Kim et al. (2017) and the third crossover from interlayer intercalation regime to growth of ordered alkali metal layer on the surface.
II.2 Holstein polarons in “surface doping” regime
The most intriguing effects in the electronic structure are observed in the “surface doping” regime, corresponding to dosing times between first and second crossover. Figure 3(a,b) shows the electronic band structure of MoSe2 at Fermi level at 120 s Rb-dosing on the surface.
Figure 3: Formation of polarons in the electron-doped MoSe2: (a) Fermi surface and (b) $\Sigma$-${\rm K}$ band dispersion measured after 120 s Rb dosing; Corresponding second derivative of ARPES intensity at $\Sigma$ (c) and ${\rm K}$ (d) valleys.
The measured Fermi surface shows two electron-like pockets at $\Sigma$ valley and single electron-like pocket at ${\rm K}$ valley. The second derivative images of $\Sigma$ and ${\rm K}$ valleys in Fig. 3(c,d) both show well-defined discontinuities of the spectral function separated by 20 meV. Moreover, the photoemission spectra integrated the momentum range between $\Sigma$ and ${\rm K}$ valleys in Fig. 5(a), show discrete non-dispersing energy levels with the same energy separation (dashed vertical black lines in Fig. 5(b)).
The energy of these spectral features is consistent with the calculated acoustic phonon mode, and their non-dispersing character in an extended momentum space, all point toward formation of small range polarons.Therefore, for the electron doped MoSe2 Holstein type polarons are observed at both ${\rm K}$ and $\Sigma$ valleys.
II.3 Spin-splitting at the $\Sigma$ valley in “surface doping” regime
While polarons at the spin-degenerate ${\rm K}$-valley has been observed previously for MoS2 Kang et al. (2018), polarons at the spin-polarised $\Sigma$ valleys are discovered for the first time. The spin-orbit splitting of the electron band dispersion at $\Sigma$ valley is apparent in ARPES data measured along the $\Gamma$-${\rm K}$ direction. Figure 4 shows band dispersions at the $\Sigma$ valley together with the Rashba-like model fit constrained by the band positions from the corresponding momentum distribution curves (MDCs) at the Fermi level.
Figure 4: Evolution of the spin-orbit coupling with Rb dosing: (a-e) Band dispersion at $\Sigma$ valley with Rb dosing. White doted lines are fits with Rashba-like model with $\Gamma$ been time-reversal invariant momentum (TRIM) point. (f) Extracted spin-split energy at $\Sigma$ valley.
The increase of the Fermi surface size indicates the gradual electron doping of MoSe2 with Rb dosing. From the same data one could see that the magnitude of band-splitting is also increasing with electron doping.This Rashba-like spin-splitting of the bands comes from inversion symmetry breaking due to the electric field created by electron charge transfer from dosed alkali atoms at the surface. Such inversion symmetry breaking on the surface has been seen at the valence band of semiconducting Kang et al. (2017); Kim et al. (2017) as well as at the conduction band of semimetallic Clark et al. (2019) alkali dosed TMDC materials. Figure 4(f) shows how the magnitude of the observed splitting consistently and monotonically becomes larger as the surface charge increases. This observation aligns perfectly with our understanding that the band splitting is Rashba-like and arises from the combination of spin-orbit coupling with the increasing inversion symmetry-breaking electric field at the surface.
II.4 Plasmonic flat states in intercalation regime
At dosing times above 340 s one observes a formation of the flat states in the vicinity of the CBM with much higher binding energies than for Holstein polaronic modes (white arrow at Fig. 5(a)).
Figure 5: Crossover from polarons to plasmons in the MoSe2 electronic structure with Rb dosing: (a) Wide energy range $\Gamma$-${\rm K}$ band dispersion measured after 340 s Rb dosing. A white arrow indicates a flat state between $\Sigma$ and ${\rm K}$ valleys; (b) Integrated Energy Distribution Curves (EDCs) between $\Sigma$ and ${\rm K}$ valleys (yellow box in (a)).
To emphasize the difference between these new flat states and Holstein polarons, we analysed momentum integrated EDCs between $\Sigma$ and ${\rm K}$ valleys in Fig.5(a). Fig. 5(b) shows that indeed two different features are visible in the integrated EDCs. The characteristic binding energies of the polaronic states in the “surface doping” regime are constant and correspond to the acoustic phonon frequency $\simeq$20 meV, but the higher binding energies of the flat states in the “interlayer intercalation” regime significantly increase with dosing. While binding energies of these flat states are an order of magnitude larger that of Holstein polarons, they are also an order of magnitude smaller then the 3.4 eV bulk plasmon frequency of metallic Rb Vom Felde et al. (1987). Binding energies of the flat states as a function 2D electron carrier density can be fitted with plasmon-like dispersion $\hbar\omega_{p}=\sqrt{(n_{e}^{\Sigma}+n_{e}^{K})e^{2}/m\epsilon}$ The obtained dielectric constant $\epsilon$$\simeq$14.3 is comparable with the calculated in-plane dielectric constant for monolayer MoSe2 Laturia et al. (2018). This suggests that these flat states may indeed be plasmonic polarons, similar to the case of EuO Riley et al. (2018) and anatase TiO2 Ma et al. (2020). Plasmonic polarons have been observed by ARPES in electron doped MoS2 Caruso et al. (2021), MoS2/TiO2 Xiang et al. (2023), HfS2 Emeis et al. (2023) and WS2 Ulstrup et al. (2023).
III Conclusion
By studying electronic structure evolution of MoSe2 with Rb surface dosing we have identified three different regimes: surface doping, inter-layer intercalation and ordered alkali metal layer formation. Figure 6(a) shows the two dimensional (2D) electron carrier density as a function of the alkali metal dosing time, calculated using the Luttinger’s theorem and experimentally determined sizes of the Fermi surface. In the “surface doping” regime we observe continuous electron doping of MoSe2 with corresponding increase of the Fermi surface size and value of the Rashba-like spin-orbit splitting.
Figure 6: Phase diagram and Fermi surface of MoSe2 with electron doping: (a) 2D electron carrier density as a function of the Rb-dosing time; (b) Measured Fermi surface of the electron doped MoSe2 after 140 s Rb dosing. (c) Fermi surface extrapolated for maximum Tc $\sim$7 K carrier density doping level Shi et al. (2015); (d,e, f) Schematic of the crystal structure for surface dosing, intercalation and metallic Rb regimes. L1 and L2 notes topmost and second topmost layers, correspondingly.
The changing nature of the quasiparticles at the Fermi level due to electron-phonon coupling, as well as the evolution of the Fermi surface topology and the electronic density of states, plays a crucial role in understanding superconductivity in electron-doped TMDC semiconductors discovered using liquid gating devices Taniguchi et al. (2012); Ye et al. (2012); Shi et al. (2015); Lu et al. (2015). In the case of MoSe2 the superconductivity occurs at high 2D electron carrier densities $\geq$8$\cdot$1013 cm-2 and maximum Tc $\sim$7 K is achieved at electron density $\simeq$16.9$\cdot$1013 cm-2, correspondingly Shi et al. (2015).
While our results show that the increasing surface electric field description is no longer applicable when we dose Rb atoms above 120 s, we could extrapolate a gradual increase of the 2D electron carrier density with the alkali metal dosing time within the ”surface doping” regime up to the superconducting transition. In this case, using our ARPES data we discover that maximum Tc doping level corresponds to the Fermi surface topology where the electron-like pocket at ${\rm K}$ valley touches the outer electron pocket at $\Sigma$ valley (see Fig. 6(b)). In another words, the maximum Tc corresponds to the Fermi level being located exactly at the conduction band saddle point between ${\rm K}$ and $\Sigma$ valleys. This can lead to a dramatic enhancement of pairing strength, due to strong increase of the momentum resolved electron-phonon coupling constant $\lambda$, like in case of Rb intercalated PtTe2  Wu et al. (2021). It can be assumed that with a further increase in doping, the electronic structure will undergo a Lifshitz transition, significantly reducing the density of states at the Fermi level and, correspondingly, suppressing electronic pairing. This can naturally explain the observed superconducting dome for Tc vs. doping phase diagram Ye et al. (2012); Shi et al. (2015).
Using ionic liquid gating devices, it was found that for electric-field-induced superconductivity in MoS2 the observed upper critical field is four-five times exceeds the Pauli limit Lu et al. (2015); Saito et al. (2016). This violation of the paramagnetic limit and significant spin-orbit coupling at the Fermi level suggest a non-spin-singlet pairing that is robust against an external magnetic field. Ising pairing for the valley alternating spin-split Fermi surface has been suggested in a several theoretical works for non-centrosymmetric TMDC superconductors Zhou et al. (2016); Ilić et al. (2017); Wang et al. (2019); Liu et al. (2020); Wickramaratne et al. (2020); Semenov (2024). However, all these models consider basic Fermi surface topology with spin-degenerate states at ${\rm K}$ valleys only. At the same time our study shows the direct evidence of Rashba-like splitting of the bands at $\Sigma$ valley for such in-situ electron doped materials.
Important consequence of this discovery is that the significant part of the total Fermi surface will be spin split at the optimal doping. This agrees with the theoretical prediction Zhao et al. (2020) that with anti-symmetric field effect doping the resulting Fermi surface is dominantly formed by the pockets around $\Sigma$ valleys (only $\sim$30of its area is around ${\rm K}$ valleys as shown in Fig.6(b)). Thus any realistic pairing model of Ising superconductivity in electron doped TMDCs should include such composite spin-split multi-valley Fermi surface topology and corresponding interband scattering processes.
To summarise, we observed spin-polarised Holstein type polaronic quasiparticles at Rb dosed MoSe2 surfaces. This is a direct evidence of the predicted Rashba-like spin-orbit split polarons Covaci and Berciu (2009). We found that beyond the ”surface doping” regime threshold, when dosed alkali atoms intercalate into the van der Waals (vdW) gap between the layers, new plasmon-like quasiparticles emerge. Our findings not only suggest the crucial role of polaron formation for the superconductivity in electron-doped transition-metal dichalcogenides, but also provide experimental foundation for theoretical models of Ising-type superconductivity in these materials.
IV Methods
High quality single crystals of MoSe2 were grown by the chemical vapor transport technique Legma et al. (1993). ARPES measurements were performed at the I05 beamline at the Diamond Light Source, UK Hoesch et al. (2017). The photoelectron’s energy and angular distributions were analysed with a SCIENTA R4000 hemispherical analyser. The angular resolution was 0.2, and the overall energy resolution was better than 5 meV. P-polarised 60 eV photons corresponding to the $\Gamma$ point of the bulk Brillouin zone were used for high-resolution ARPES, and 120 eV photons for Se 3d core level measurements. Single crystal samples were cleaved and measured in ultra high-vacuum at temperatures below 10 K and pressures below 1$\cdot$10-10 mbar. Rb atoms were dosed in several sequences on a cold MoSe2 surface using a commercial SAES alkali-metal dispenser. The density functional theory (DFT) calculations including spin-orbit coupling with generalized gradient approximation (GGA) were performed using WEIN2K package Blaha et al. (2020). Spectral function simulation for polarons has been done by using momentum average (MA) approximation Goodvin et al. (2006) with 20 meV acoustic phonons.
V Acknowledgements
We acknowledge Diamond Light Source for time on Beamline I05-ARPES under Proposal NT26631. S. W. Jung was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1C1C1004978) and supported by Learning & Academic research institution for Master’s·PhD students, and Postdocs (LAMP) Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2023-00301974).
Supplementary
Density Functional Theory (DFT) calculations
Band structure DFT calculations including spin-orbit coupling show that bottom of the conduction band at $\Sigma$-valley is lower for bulk MoSe2 then for bulk MoS2 and lower then bottom of the conduction band at ${\rm K}$-valley. This makes MoSe2 a much better choice to study polaron formation at $\Sigma$-valley. Importantly, in case of broken inversion symmetry, due to electric field built up on the surface caused by the alkali metal dosing, one expect corresponding spin-splitting at $\Sigma$-valley, like in the case of monolayer MoSe2.
Figure S1: DFT+SOC calculated electronic structure of bulk MoS2, bulk MoSe2 and monolayer MoSe2.
The evolution of MoSe2 electronic structure with Rb dosing.
Photoemission data for conduction (Fig. S2) and valence (Fig. S3) bands together with corresponding Se 3d core levels data (Fig. S4) were measured for all Rb dosing sequences.
Figure S2: Conduction band electronic structure at the Fermi-level of Rb dosed MoSe2 surface.
Figure S3: Valence band electronic structure of Rb dosed MoSe2 surface.
Figure S4: Se 3p core-levels with Rb dosing. (a) Se 3d5/2/3d3/2 XPS data with Voight profile doublets fit for various dosing times. Red and blue curves indicate bulk (P1) and surface (P2) components; (b) Ratio of the bulk and surface components versus Rb dosing time; (c) Energy difference between bulk and surface components versus Rb dosing time; The background colour in (b) and (c) indicates surface dosing, intercalation and Rb metallisation regimes.
A single 3d5/2/3d3/2 doublet for a pristine MoSe2 crystal splits into two components with Rb dosing as shown in Fig. S4(a). Similar to the case of HfTe2El Youbi et al. (2020) one could fit these core level data assuming bulk (P1) and surface (P2) components due to the electron doping of the MoSe2 surface. Obtained peak area ratio and energy differences (or effective chemical shift) are shown in Fig. S4(b,c). Below critical dosing of 120 s one observes increase of the difference between bulk and surface doublets binding energies with dosing time. This corresponds to the gradual accumulation of Rb atoms, increased electron doping of the MoSe2 surface, and corresponding increase of effective electric filed.
Doping evolution of the band gap in MoSe2
Similar to the previously published ARPES data of alkali-dosed MoSe2 Kim et al. (2017); Kang et al. (2017) we observe superposition of the bulk and surface electronic states. This is especially evident at top of valence band at ${\rm K}$-valley where one detects weak surface related features at lower binding energies than bulk ones. At higher doping levels this might lead to a shift of the valence band maxima (VBM) from $\Gamma$ to ${\rm K}$ valley and corresponding transition from indirect to direct band gap electronic structure. This has been clearly observed by ARPES in the case of MoTe2 Kang et al. (2017). However, in the case of MoSe2 this is less obvious from the ARPES data Kim et al. (2017); Kang et al. (2017).
In Fig. S5(b,c) photoemission peak maxima of the conduction band for both $\Sigma$ to ${\rm K}$ valleys shift towards higher binding energies, away from the Fermi level. However withing experimental broadening of corresponding spectra, one could not with absolute certainty conclude, that conduction band minima is shifting from $\Sigma$ to ${\rm K}$ valley. Therefore, despite obvious spectral intensity shift from $\Sigma$ to ${\rm K}$ valley in Fig. S2, we could not presume the crossover to a monolayer conduction band dispersion in alkali-metal dosed MoSe2. Correspondingly, in Fig. S5(e,f) photoemission peak maxima of the valence band for both $\Gamma$ to ${\rm K}$ valleys shift towards Fermi level with dosing. Withing experimental broadening of corresponding spectra, one could not determine, that valence band maxima is shifting from $\Gamma$ to ${\rm K}$ valley. The lower intensity surface related features might be especially well masked by the bulk component at $\Gamma$ valley as compared to ${\rm K}$ valley as seen in Fig. S7. Additional broadening of the spectra due to significant kz dispersion at the centre of the Brillouin zone make distinguishing bulk and surface components there even more challenging. We can not exclude electron doping induced indirect to direct band gap transition in MoSe2, however this might be beyond ”surface doping” regime achievable with alkali metal dosing at the surface.
Figure S5: Evolution of the conduction band minima and valence band maxima with Rb dosing: (a) Conduction bands dispersion along $\Gamma$-${\rm K}$ direction; (b) Dispersion of conduction band minima with surface doping obtained from EDC at $\Sigma$ point; (c) Dispersion of conduction band minima with surface doping obtained from EDC at ${\rm K}$ point; (d) Valence bands dispersion along $\Gamma$-${\rm K}$ direction. L1 and L2 notes topmost and second topmost layers, correspondingly; (e) Dispersion of Valence band maxima with surface doping obtained from EDC at $\Sigma$ point; (f) Dispersion of Valence band maxima with surface doping obtained from EDC at ${\rm K}$ point;
Figure S6: Evolution of conduction band minima with Rb dosing: (a) Energy distribution curves (EDCs) at $\Sigma$ point; (b) Energy distribution curves (EDCs) at ${\rm K}$ point.
Figure S7: Evolution of valence band maxima with Rb dosing: (a) Energy distribution curves (EDCs) with fits at $\Gamma$ point; (b) Energy distribution curves (EDCs) with fits at ${\rm K}$ point.
Figure S8: Constant energy maps (hv=180eV) of MoSe2 surface after 140 s Rb-dosing:
(a) Momentum distribution map at Fermi level and (b) Momentum distribution map at the top of the valence band.
Rashba-like splitting at $\Sigma$ valley
To analyse spin-splitting at $\Sigma$ valley as a function of surface doping we used Rashba-like model with $\Gamma$ as time-reversal invariant momenta (TRIM) point shown in Fig. S10(a).
The effective Hamiltonian can be written as:
$H(\vec{k})=\frac{\hbar^{2}}{2m^{*}}(\vec{k}-\vec{k_{0}})^{2}\pm\alpha_{R}(\hat{n}\times\vec{\sigma})\cdot\vec{k}+E_{offset}$ (1)
where k0 is the momentum of the conduction band minimum near $\Sigma$ without spin-orbit coupling, $\sigma$ is Pauli matrices, $E_{offset}$ is the chemical shift due to the electron doping via Rb dosing and m is the effective mass.
From equation (1), one can make the system of equations with four $k_{F}$ (S1,S2,S3 and S4) shown in Fig. S10.
$\frac{\hbar^{2}}{2m^{*}}(k_{S1}-k_{0})^{2}-\alpha_{R}k_{S1}+E_{offset}=0$ (2)
$\frac{\hbar^{2}}{2m^{*}}(k_{S2}-k_{0})^{2}+\alpha_{R}k_{S2}+E_{offset}=0$ (3)
$\frac{\hbar^{2}}{2m^{*}}(k_{S3}-k_{0})^{2}+\alpha_{R}k_{S3}+E_{offset}=0$ (4)
$\frac{\hbar^{2}}{2m^{*}}(k_{S4}-k_{0})^{2}-\alpha_{R}k_{S4}+E_{offset}=0$ (5)
By subtracting the equation (5) with the equation (2), one can make
$\frac{\hbar^{2}}{2m^{*}}(k_{S1}+k_{S4}-2k_{0})-\alpha_{R}=0$ (6)
Likewise, one can subtract equation (4) from equation (3).
$\frac{\hbar^{2}}{2m^{*}}(k_{S2}+k_{S3}-2k_{0})+\alpha_{R}=0$ (7)
From equation (6) and (7), one can derive $k_{0}$ from four $k_{F}$ values
$k_{0}=\frac{k_{S1}+k_{S2}+k_{S3}+k_{S4}}{4}$ (8)
By replacing $k_{0}$ in equation (6) and (7), one can derive effective Rashba parameter $\alpha_{R}$
$\alpha_{R}=\frac{\hbar^{2}}{2m^{*}}\frac{(k_{S1}+k_{S4}-k_{S2}-k_{S3})}{2}$ (9)
From equations (1), (8) and (9), the splitting energy $E_{splitting}$ can be derived as
$E_{splitting}=2\alpha_{R}k_{0}-\frac{2m^{*}\alpha_{R}^{2}}{\hbar^{2}}$ (10)
To fit ARPES data, we used constant DFT value for the effective mass $m^{*}=0.54m_{e}$ at $\Sigma$ valley and $k_{F}$ values obtained from MDC fits shown in Fig. S9.
Figure S9: Momentum distribution curves (MDCs) of $\Sigma$ valleys at Fermi-level.
The effective Rashba parameter $\alpha_{R}$ versus Rb dosing time is shown in Fig. S10(b). Corresponding extracted spin-split energy $E_{splitting}$ at $\Sigma$ valley as a function of Rb dosing time is shown in the main text Fig. 4.
Figure S10: Spin-splitting at $\Sigma$ valley: (a) Rashba-like model with $\Gamma$ as time-reversal invariant momentum (TRIM) point; (b) Extracted Rashba parameter ($\alpha_{R}$).
For comparison, for Au(111) and Cu(111) surface states measured Rashba parameter $\alpha_{R}$ = 330 and 38 meVÅ give energy splitting $E_{splitting}$ = 110 and 16 meV, correspondingly.

### v2

@@
 The resulting polaronic bound state can be described by Fröhlich or Holstein Hamiltonians.
-The Fröhlich model describes the interaction of electrons with optical phonons, which leads to the formation of large polarons, whose radii are much larger than the lattice constant.
-The Holstein model, on the other hand, describes the interaction with acoustic phonons, which leads to the formation of small polarons, whose radii are of the same order of magnitude as the lattice constant Emin (2013).
+The Fröhlich model describes the interaction of electrons with phonons, which leads to the formation of large polarons, whose radii are much larger than the lattice constant.
+The Holstein model, on the other hand, describes the short-range interaction with phonons, which typically leads to the formation of small polarons, whose radii are of the same order of magnitude as the lattice constant Emin (2013).
 Formation of the polarons due to a momentum dependent electron-phonon coupling leads to new features in the electronic spectral function at specific electron valleys, e.g.
@@
 The electron scattering vectors $q_{1}$ and $q_{2}$ correspond to either $\Gamma$-${\rm K}$ or $\Gamma$-${\rm M}$ phonon wave vectors.
-For both MoS2 and MoSe2 the acoustic phonon modes have similar energies for ${\rm K}$ and ${\rm M}$ and the dominant contribution to electron-phonon coupling arises from the highest energy acoustic phonon mode.
+The dominant contribution to electron-phonon coupling arises from the highest energy acoustic phonon mode.
 This behaviour has been demonstrated both experimentally Zeng et al.
@@
 (2012) for MoS2.
+As in the case of MoS2 Kaasbjerg et al.
+(2012); Zhao et al.
+(2018), for MoSe2 the acoustic phonon modes have essentially flat dispersion over an extended region of $k$-space between the ${\rm M}$ and ${\rm K}$ points Bae et al.
+(2022).
+Thus, despite their acoustic nature, these phonon modes with characteristic frequency $\Omega_{ph}$ define the energy scale for electron-phonon coupling and corresponding polaron formation Zhao et al.
+(2018); Bae et al.
+(2022).
 However, only ${\rm K}$ valley Holstein polarons (Fig.
@@
 (2021).
-Figure 2: Evolution of the MoSe2 electronic structure with Rb dosing: (a-a5) ARPES spectra of $\Gamma$-${\rm K}$ conduction band dispersions measured near the Fermi energy for dosing 0, 20, 60, 100, 220, 580s.
+Figure 2: Evolution of the MoSe2 electronic structure with Rb dosing: (a-a5) ARPES spectra of $\Gamma$-${\rm K}$ conduction band dispersions measured near the Fermi energy for cumulative dosing time 0, 20, 60, 100, 220, 580s.
 (b-b5) Corresponding valence band spectra in a wide energy range.
@@
 S1).
-For the alkali-metal dosing on MoSe2 surface one expects to have larger SOC band-splitting as in the case of MoS2.
+For the alkali-metal dosing on MoSe2 surface one expects to have larger SOC band-splitting than in the case of MoS2.
 The accessibility of the $\Sigma$ valley, as well as the ${\rm K}$ valley expected at higher doping Kim et al.
-(2016, 2017), makes MoSe2 the best candidate to study physics of multi-valley polaron modes using high-resolution ARPES.
+(2016, 2017), makes MoSe2 the best candidate to study the physics of multi-valley polaron modes using high-resolution ARPES.
 II Results and discussion
@@
 After initial dosing the maximum spectral weight at the Fermi level is observed around $\Sigma$ valley, showing an indirect band gap $E_{g}^{\Gamma-\Sigma}$ $\sim$1.4 eV.
-However, with increasing amount of dosed Rb the relative intensity of $\Sigma$ valley decreases, whereas the intensity of ${\rm K}$ valley increases.
-Correspondingly, it appears that the conduction band minimum shifts from $\Sigma$ to ${\rm K}$ valley as in case of monolayer MoSe2 (see Fig.
+However, with increasing amount of dosed Rb the relative intensity of the $\Sigma$ valley decreases, whereas the intensity of ${\rm K}$ valley increases.
+Correspondingly, it appears that the conduction band minimum shifts from $\Sigma$ to ${\rm K}$ valley as in the case of monolayer MoSe2 (see Fig.
 2(a1-a3)).
 Withing observed energy broadening the band minima at $\Sigma$ and ${\rm K}$ are very close (see Supplementary Fig.
-S5), incidentally making electron doped MoSe2 perfect system for studying multi-valley physics.
+S5), incidentally making electron doped MoSe2 a perfect system for studying multi-valley physics.
 Simultaneously with the shift of the CBM intensity, the intensity of the valence band compared to the conduction band decreases (see Fig.
@@
 3(c,d) both show well-defined discontinuities of the spectral function separated by 20 meV.
-Moreover, the photoemission spectra integrated the momentum range between $\Sigma$ and ${\rm K}$ valleys in Fig.
+Moreover, the photoemission spectra integrated withing a selected momentum range between $\Sigma$ and ${\rm K}$ valleys in Fig.
 5(a), show discrete non-dispersing energy levels with the same energy separation (dashed vertical black lines in Fig.
@@
 (2018), polarons at the spin-polarised $\Sigma$ valleys are discovered for the first time.
-The spin-orbit splitting of the electron band dispersion at $\Sigma$ valley is apparent in ARPES data measured along the $\Gamma$-${\rm K}$ direction.
+The spin-orbit splitting of the electron band dispersion at the $\Sigma$ valley is apparent in ARPES data measured along the $\Gamma$-${\rm K}$ direction.
+This agrees with theoretical predictions of lifting the spin degeneracy under electron doping and much larger spin-orbit splitting at $\Sigma$ compared to ${\rm K}$ due to distinct orbital character of conduction bands at these valleys Brumme et al.
+(2015); Clark et al.
+(2019); Boccuni et al.
+(2024).
 Figure 4 shows band dispersions at the $\Sigma$ valley together with the Rashba-like model fit constrained by the band positions from the corresponding momentum distribution curves (MDCs) at the Fermi level.
@@
 The increase of the Fermi surface size indicates the gradual electron doping of MoSe2 with Rb dosing.
-From the same data one could see that the magnitude of band-splitting is also increasing with electron doping.This Rashba-like spin-splitting of the bands comes from inversion symmetry breaking due to the electric field created by electron charge transfer from dosed alkali atoms at the surface.
+From the same data one could see that the magnitude of band-splitting is also increasing with electron doping.
+This Rashba-like spin-splitting of the bands comes from inversion symmetry breaking due to the electric field created by electron charge transfer from dosed alkali atoms at the surface.
 Such inversion symmetry breaking on the surface has been seen at the valence band of semiconducting Kang et al.
@@
 This observation aligns perfectly with our understanding that the band splitting is Rashba-like and arises from the combination of spin-orbit coupling with the increasing inversion symmetry-breaking electric field at the surface.
+Moreover, the constancy of intensity for the two components with dosing rules out the case of superposition of the surface and subsurface as a possible origin of the observed splitting.
 II.4 Plasmonic flat states in intercalation regime
@@
 In the “surface doping” regime we observe continuous electron doping of MoSe2 with corresponding increase of the Fermi surface size and value of the Rashba-like spin-orbit splitting.
+In this regime, before interlayer intercalation occurs, Rb-dosing on the surface is comparable to doping induced by the electrostatic gating using ionic liquid.
+Therefore, the electronic structure on MoSe2 under Rb dosing in “surface doping” regime corresponds to the electronic structure of MoSe2 in ionic liquid gating device under positive gate voltage (electron doping).
 Figure 6: Phase diagram and Fermi surface of MoSe2 with electron doping: (a) 2D electron carrier density as a function of the Rb-dosing time; (b) Measured Fermi surface of the electron doped MoSe2 after 140 s Rb dosing.
@@
 While our results show that the increasing surface electric field description is no longer applicable when we dose Rb atoms above 120 s, we could extrapolate a gradual increase of the 2D electron carrier density with the alkali metal dosing time within the ”surface doping” regime up to the superconducting transition.
-In this case, using our ARPES data we discover that maximum Tc doping level corresponds to the Fermi surface topology where the electron-like pocket at ${\rm K}$ valley touches the outer electron pocket at $\Sigma$ valley (see Fig.
+In this case, using extrapolation from our ARPES data we conclude that maximum Tc doping level corresponds to the Fermi surface topology where the electron-like pocket at ${\rm K}$ valley touches the outer electron pocket at $\Sigma$ valley (see Fig.
 6(b)).
@@
 At the same time our study shows the direct evidence of Rashba-like splitting of the bands at $\Sigma$ valley for such in-situ electron doped materials.
-Important consequence of this discovery is that the significant part of the total Fermi surface will be spin split at the optimal doping.
+An important consequence of this discovery is that a significant part of the total Fermi surface will be spin split at the optimal doping.
 This agrees with the theoretical prediction Zhao et al.
@@
 Thus any realistic pairing model of Ising superconductivity in electron doped TMDCs should include such composite spin-split multi-valley Fermi surface topology and corresponding interband scattering processes.
-To summarise, we observed spin-polarised Holstein type polaronic quasiparticles at Rb dosed MoSe2 surfaces.
-This is a direct evidence of the predicted Rashba-like spin-orbit split polarons Covaci and Berciu (2009).
-We found that beyond the ”surface doping” regime threshold, when dosed alkali atoms intercalate into the van der Waals (vdW) gap between the layers, new plasmon-like quasiparticles emerge.
-Our findings not only suggest the crucial role of polaron formation for the superconductivity in electron-doped transition-metal dichalcogenides, but also provide experimental foundation for theoretical models of Ising-type superconductivity in these materials.
+To summarise, we observed, for the first time, spin-polarised Holstein type polaronic quasiparticles on Rb dosed MoSe2 surfaces.
+This is direct evidence of the predicted Rashba-like spin-orbit split polarons Covaci and Berciu (2009).
+We found that beyond the “surface doping” regime threshold, when dosed alkali atoms intercalate into the van der Waals (vdW) gap between the layers, completely new plasmon-like quasiparticles emerge.
+Our findings not only suggest the crucial role of polaron formation for the electronic structure of electron-doped transition-metal dichalcogenides, but also provide experimental foundation for theoretical models of Ising-type superconductivity in these materials.
 IV Methods
@@
 Rb atoms were dosed in several sequences on a cold MoSe2 surface using a commercial SAES alkali-metal dispenser.
-The density functional theory (DFT) calculations including spin-orbit coupling with generalized gradient approximation (GGA) were performed using WEIN2K package Blaha et al.
+The density functional theory (DFT) calculations including spin-orbit coupling with generalized gradient approximation (GGA) were performed using Wien2K package Blaha et al.
 (2020).
 Spectral function simulation for polarons has been done by using momentum average (MA) approximation Goodvin et al.
-(2006) with 20 meV acoustic phonons.
+(2006) with 20 meV constant phonon energy.
 V Acknowledgements
@@
 Below critical dosing of 120 s one observes increase of the difference between bulk and surface doublets binding energies with dosing time.
-This corresponds to the gradual accumulation of Rb atoms, increased electron doping of the MoSe2 surface, and corresponding increase of effective electric filed.
+This corresponds to the gradual accumulation of Rb atoms, increased electron doping of the MoSe2 surface, and corresponding increase of effective electric field.
+Beyond this critical dosing time, Rb atoms penetrate the van der Waals gap underneath the top layer, thereby physically and electronically isolating the topmost MoSe2 layer.
+This clearly marks the second crossover from surface doping to the interlayer intercalation regime.
 Doping evolution of the band gap in MoSe2
@@
 To analyse spin-splitting at $\Sigma$ valley as a function of surface doping we used Rashba-like model with $\Gamma$ as time-reversal invariant momenta (TRIM) point shown in Fig. S9(a).
+Figure S9: Spin-splitting at $\Sigma$ valley: (a) Rashba-like model with $\Gamma$ as time-reversal invariant momentum (TRIM) point; (b) Extracted Rashba parameter ($\alpha_{R}$).
 The effective Hamiltonian can be written as:
@@
 $E_{splitting}=2\alpha_{R}k_{0}-\frac{2m^{*}\alpha_{R}^{2}}{\hbar^{2}}$ (10)
-To fit ARPES data, we used constant DFT value for the effective mass $m^{*}=0.54m_{e}$ at $\Sigma$ valley and $k_{F}$ values obtained from MDC fits shown in Fig. S9.
+To fit the ARPES data, we used constant DFT value for the effective mass $m^{*}=0.54m_{e}$ at $\Sigma$ valley and $k_{F}$ values obtained from MDC fits shown in Fig.
+S10.
+From these $k_{F}$ values we estimate the experimental asymmetry of the Rashba splitting at $\Sigma$ valley of about $\sim$9%.
 Figure S10: Momentum distribution curves (MDCs) of $\Sigma$ valleys at Fermi-level.
 The effective Rashba parameter $\alpha_{R}$ versus Rb dosing time is shown in Fig.
-S10(b).
+S9(b).
 Corresponding extracted spin-split energy $E_{splitting}$ at $\Sigma$ valley as a function of Rb dosing time is shown in the main text Fig.
 4.
-Figure S10: Spin-splitting at $\Sigma$ valley: (a) Rashba-like model with $\Gamma$ as time-reversal invariant momentum (TRIM) point; (b) Extracted Rashba parameter ($\alpha_{R}$).
 For comparison, for Au(111) and Cu(111) surface states measured Rashba parameter $\alpha_{R}$ = 330 and 38 meVÅ give energy splitting $E_{splitting}$ = 110 and 16 meV, correspondingly.